# Retrieval-Augmented Generation with Estimation of Source Reliability

### v2

@@ Abstract
Retrieval-Augmented Generation (RAG) is an effective approach to enhance the factual accuracy of large language models (LLMs) by retrieving information from external databases, which are typically composed of diverse sources, to supplement the limited internal knowledge of LLMs. However, the standard RAG often risks retrieving incorrect information, as it relies solely on relevance between a query and a document, overlooking the heterogeneous reliability of these sources. To address this issue, we propose Reliability-Aware RAG (RA-RAG), a multi-source RAG framework that estimates the reliability of sources and leverages this information to prioritize highly reliable and relevant documents, ensuring more robust and accurate response generation. Specifically, RA-RAG first estimates source reliability by cross-checking information across multiple sources. It then retrieves documents from the top-$\kappa$ reliable and relevant sources and aggregates their information using weighted majority voting (WMV), where the selective retrieval ensures scalability while not compromising the performance. Comprehensive experiments show that RA-RAG consistently outperforms baselines in scenarios with heterogeneous source reliability while scaling efficiently as the number of sources increases. Furthermore, we demonstrate the ability of RA-RAG to estimate real-world sources’ reliability, highlighting its practical applicability.
Machine Learning, ICML
@@ 1 Introduction
Large language models (LLMs) have demonstrated remarkable performance across various tasks (Zhao et al., 2023b; Brown et al., 2020). However, they often produce incorrect outputs, particularly when handling up-to-date knowledge that is absent from their internal knowledge (Shuster et al., 2021; Zhang et al., 2024; Dhuliawala et al., 2024; Huang et al., 2023; Zhao et al., 2023a). To address this limitation, retrieval-augmented generation (RAG) (Guu et al., 2020; Lewis et al., 2020; Asai et al., 2024; Yan et al., 2024) has emerged as a promising approach, leveraging external knowledge from large-scale databases that integrate an extensive set of sources (Vu et al., 2024; Kasai et al., 2024).
@@
Although external databases provide valuable information, they also risk retrieving incorrect information from unreliable sources (Pan et al., 2023; Chen et al., 2024; Greshake et al., 2023). This vulnerability stems from a fundamental limitation of retrieval, which relies solely on relevance measures between queries and documents (Robertson & Walker, 1994; Karpukhin et al., 2020; Ni et al., 2022; Izacard et al., 2022), overlooking source reliability heterogeneity. Furthermore, malicious sources can exploit this limitation by crafting highly relevant yet incorrect documents, leading to misleading outputs (Zhong et al., 2023; Zou et al., 2024). While existing methods (Weller et al., 2024; Xiang et al., 2024; Deng et al., 2024; Pan et al., 2024) attempt to mitigate this issue by refining retrieved documents, they do not address the retrieval problem itself, allowing unreliable sources to dominate the retrieval process.
@@
To overcome these challenges, we propose Reliability-Aware RAG (RA-RAG), a new multi-source RAG framework that estimates source reliability and effectively integrates it into both the retrieval and aggregation processes. Compared to standard RAG, which retrieves documents without distinguishing between sources, RA-RAG performs source-level retrieval and aggregates information based on estimated reliability using weighted majority voting (WMV), as illustrated in Figure 1. Specifically, RA-RAG consists of two steps. First, given a set of fact-checking queries, we estimate source reliability by cross-checking information across multiple sources without requiring manual fact-checking. This is achieved by leveraging RAG’s ability to automatically retrieve and generate responses (Section 4.2). Second, using the estimated reliability, we propose $\kappa$-reliable and relevant source selection ($\kappa$-RRSS) for WMV, where RA-RAG consults only a small number of reliable sources with relevant documents. This enhances robustness against unreliable sources while maintaining computational scalability without compromising performance.
The effectiveness of RA-RAG stems from its ability to estimate source reliability, a crucial first step in combating misinformation (Popat et al., 2017; Baly et al., 2018, 2020; Burdisso et al., 2024). While source reliability remains underexplored in RAG despite its significance, RA-RAG explicitly incorporates it to improve retrieval and answer generation. Comprehensive experiments and analyses demonstrate that RA-RAG not only effectively estimates source reliability but also robustly aggregates information from multiple sources with heterogeneous reliability. Moreover, it remains scalable even as the number of sources increases. Furthermore, our method effectively estimates the reliability of real-world sources, highlighting its practical applicability.
Our main contributions are summarized as follows:
  • We propose RA-RAG, a multi-source RAG framework that estimates source reliability by cross-checking information across sources without relying on manual fact-checking (Section 4.2). Based on the estimated reliability, it retrieves reliable and relevant documents by $\kappa$-RRSS and aggregates them with WMV, generating robust answers while remaining scalable to a large number of sources (Section 4.1).
@@ 2 Related Works
Retrieval-augmented generation. RAG enhances LLM performance by retrieving information from external databases, reducing hallucinations, and improving access to up-to-date knowledge (Lewis et al., 2020; Guu et al., 2020). Since irrelevant documents are prevalent in retrieval results, many studies have focused on enhancing RAG’s robustness through advanced retrieval methods, such as adaptive retrieval (Asai et al., 2024; Jiang et al., 2023), reranking retrieved documents (Glass et al., 2022), and query reformulation (Wang et al., 2023; Ma et al., 2023). While these approaches improve the retrieval process, they still rely on relevance measures between queries and documents, leaving them vulnerable to misinformation. (Zou et al., 2024).
Robust RAG against misinformation. In response to misinformation risks in RAG, several robust methods have been proposed, primarily focusing on improving answer generation after retrieval. Weller et al. (2024); Xiang et al. (2024) utilize majority voting, which is effective only when most retrieved documents are trustworthy. Deng et al. (2024) evaluates document credibility using LLMs’ internal knowledge, but this approach is inherently limited as it misaligns with RAG’s core rationale of leveraging external knowledge to address LLMs’ limitations. Pan et al. (2024) assigns binary credibility scores (high/low) to retrieved documents based on source reputation and incorporates them into prompts. However, this approach is unsuitable for social media, where reputation is obscure and manipulable, and it requires fine-tuning an LLM. In contrast, to the best of our knowledge, RA-RAG is the first approach to integrate source reliability estimation within the RAG framework.
@@ 3 Problem Formulation
In this section, we first introduce the standard RAG framework in Section 3.1 which has been widely used in previous works but has a clear limitation: overlooking the source reliability heterogeneity. To address this, we introduce a multi-source RAG framework that accounts for source reliability in Section 3.2, followed by a discussion of its key challenges.
@@ 3.2 Multi-source RAG with source reliability
We introduce a multi-source RAG framework that explicitly distinguishes between the sources of documents and incorporates source reliability. Let $N$ be the number of distinct sources contributing to the database $\mathcal{D}$. We partition the database as $\mathcal{D}=\bigcup_{i=1}^{N}\mathcal{S}_{i}$, where $\mathcal{S}_{i}$ is the set of documents originating from source $i\in[N]$. Such a partition of dataset $\mathcal{D}$ enables the system to account for the reliability of each document’s source, based on weighted majority voting (WMV). For a given query $q$, let $\hat{y}_{i}=\mathcal{G}(q,\mathcal{R}(q,\mathcal{S}_{i}))$ represent the generated response using retrieved documents exclusively from source $S_{i}$. Once the probability that a retrieved document from source $i$ is correct is estimated as $v_{i}$, and a set of candidate responses $\mathcal{M}$ is obtained from $\hat{y}_{i}$’s, we apply WMV to aggregate the responses as follows:
$\vspace{.5em}\hat{y}=\operatorname*{arg\,max}_{u\in\mathcal{M}}\sum_{i\in[N]}v_{i}\mathbbm{1}(\hat{y}_{i}=u)\;.$
@@
Inherent issues with LLMs. LLMs may generate hallucinations or misaligned answers influenced by their internal knowledge (Kaddour et al., 2023; Ji et al., 2023; Xie et al., 2024; Kortukov et al., 2024; Xu et al., 2024), distorting their alignment with retrieved documents and complicating the WMV process. Additionally, LLMs often generate semantically identical responses with paraphrasing, making response aggregation of $\hat{y}_{i}$’s more challenging.
@@ 4 Method: RA-RAG
We propose Reliability-Aware RAG (RA-RAG) to address key challenges in multi-source RAG. RA-RAG first estimates source reliability using an iterative reliability estimation algorithm, which cross-checks information across multiple sources through fact-checking queries designed to verify documents within each source (Section 4.2). Leveraging RAG’s ability to retrieve relevant documents and generate responses automatically, RA-RAG enables automated reliability estimation without manual fact-checking. It then aggregates responses from different sources based on the estimated reliability (Section 4.1).
For ease of presentation, we first introduce the aggregation process in Section 4.1, then propose the iterative reliability estimation method in Section 4.2. We also provide an illustrative overview of RA-RAG in Appendix A.
@@ 4.1 Aggregation process
Although the instruction prompt guides the model to output “I don’t know” (IDK) when there is no relevant information, LLMs may still produce misaligned responses, undermining effective aggregation. To address this, a filtering function $f_{\text{align}}$ is necessary to detect and replace misaligned responses with IDK. In this work, we utilize AlignScore (Zha et al., 2023), which evaluates the factual consistency of a response $\hat{y}_{i}$ relative to the query $q$ and retrieved documents $\mathcal{R}(q,\mathcal{S}_{i})$:
$\vspace{.5em}f_{\text{align}}(\hat{y}_{i},q,\mathcal{R}(q,\mathcal{S}_{i}))=\begin{cases}\text{IDK}&\text{if~{}}\mathcal{E}(\hat{y}_{i};q,\mathcal{R}(q,\mathcal{S}_{i}))<\tau\\ \hat{y}_{i}&\text{otherwise}\end{cases},$
where $\mathcal{E}$ represents AlignScore function and $\tau$ is threshold. For simplicity, we omit $\mathcal{E}$ in $f_{\text{align}}(\hat{y}_{i},q,\mathcal{R}(q,\mathcal{S}_{i}))$. Further details on the filtering method and threshold are provided in Appendix B. By applying this filtering method, we obtain a refined set of candidate responses:
@@
Additionally, since LLMs often paraphrase responses with equivalent meanings (e.g., “There are 24 hours in a day.” vs. “Each day has 24 hours.”), we cluster responses in $\mathcal{M}_{\text{filtered}}$ based on semantic equivalence. The refined set is denoted as $\mathcal{C}(\mathcal{M}_{\text{filtered}})$, where $\mathcal{C}$ represents a semantic clustering method. We employ the algorithm by Kuhn et al. (2023), which clusters responses that mutually entail each other using a pretrained natural language inference (NLI) model. Following Kuhn et al. (2023), we use the DeBERTa-large model (He et al., 2021) for clustering.
@@
$\vspace{.5em}\hat{y}=\!\operatorname*{arg\,max}_{u\in\mathcal{C}(\mathcal{M}_{\text{filtered}})}\sum_{i\in[N]}\!v_{i}\mathbbm{1}(f_{\text{align}}(\hat{y}_{i},q,\mathcal{R}(q,\mathcal{S}_{i}))\!=u)\;.$ (1)
To generate the final response $\hat{y}$, we select the first response in the cluster, as all responses within the cluster are considered semantically equivalent.
Efficient aggregation. In real-world applications, aggregating information from all sources can be computationally expensive, especially when the number of sources is large. To mitigate this, we propose $\kappa$-Reliable and Relevant Source Selection ($\kappa$-RRSS). This method iterates over sources in descending order of reliability and selects the first $\kappa$ sources that contain relevant information, where $\kappa<N$. A source is deemed irrelevant if its filtered response $f_{\text{align}}(\hat{y}_{i},q,\mathcal{R}(q,\mathcal{S}_{i}))$ is “I don’t know”. For the formal algorithm, please refer to Algorithm 1. Given the selected sources, denoted as $\mathcal{M}_{\kappa}$, the final response is aggregated as follows:
$\hat{y}\vspace{1em}=\!\!\!\operatorname*{arg\,max}_{u\in\mathcal{C}(\mathcal{M}_{\kappa\text{-filtered}})}\!\sum_{i\in[N]}\!v_{i}\mathbbm{1}(f_{\text{align}}(\hat{y}_{i},q,\mathcal{R}(q,\mathcal{S}_{i}))\!=u)\;,$ (2)
where $\mathcal{M}_{\kappa\text{-filtered}}$ denotes the set of responses from $\mathcal{M}_{\kappa}$ after applying $f_{\text{align}}$. By focusing on reliable and relevant sources, $\kappa$-RRSS significantly reduces inference overhead while maintaining robust performance. Furthermore, we also explore $\kappa$-Reliable Source Selection ($\kappa$-RSS), which chooses only the $\kappa$ most reliable sources without checking for relevance. Although this method further lowers computational costs, high reliability alone does not guarantee that a source contains the relevant document for a given query, potentially degrading performance. We present ablation studies in Section 5.3.
@@ 4.2 Iterative reliability estimation
To estimate source reliability and effectively aggregate outputs, we utilize the WMV method proposed by Li & Yu (2014), a simple yet effective approach for aggregating crowdsourced labels in classification tasks. Specifically, we first generate fact-checking queries for documents. For example, if a document in a source states, “COVID-19 is caused by 5G networks”, we can generate a query such as “What causes COVID-19?”. Given a set of $M$ fact-checking queries, denoted as $\{q^{j}\mid j\in[M]\}$, the iterative reliability estimation process is described as follows:

### v3

@@ Abstract
Retrieval-Augmented Generation (RAG) is an effective approach to enhance the factual accuracy of large language models (LLMs) by retrieving information from external databases, which are typically composed of diverse sources, to supplement the limited internal knowledge of LLMs. However, the standard RAG often risks retrieving incorrect information, as it relies solely on relevance between a query and a document, overlooking the heterogeneous reliability of these sources. To address this issue, we propose Reliability-Aware RAG (RA-RAG), a new multi-source RAG framework that estimates the reliability of sources and leverages this information to prioritize highly reliable and relevant documents, ensuring more robust and accurate response generation. Specifically, RA-RAG first estimates source reliability by cross-checking information across multiple sources. It then retrieves documents from the top-$\kappa$ reliable and relevant sources and aggregates their information using weighted majority voting (WMV), where the selective retrieval ensures scalability while not compromising the performance. Comprehensive experiments show that RA-RAG consistently outperforms baselines in scenarios with heterogeneous source reliability while scaling efficiently as the number of sources increases. Furthermore, we demonstrate the ability of RA-RAG to estimate real-world sources’ reliability, highlighting its practical applicability.
Retrieval-Augmented Generation with Estimation of Source Reliability
Jeongyeon Hwang1, Junyoung Park 1, Hyejin Park1, Dongwoo Kim1, Sangdon Park1, Junseul Ok1,thanks: Corresponding authors. Email: jungseul@postech.ac.kr, 1Pohang University of Science and Technology (POSTECH), South Korea
@@ 1 Introduction
Large language models (LLMs) have demonstrated remarkable performance across various tasks Zhao et al. (2023b); Brown et al. (2020). However, they often produce incorrect outputs, particularly when handling up-to-date knowledge that is absent from their internal knowledge Shuster et al. (2021); Zhang et al. (2024); Dhuliawala et al. (2024); Huang et al. (2023); Zhao et al. (2023a). To address this limitation, retrieval-augmented generation (RAG) Guu et al. (2020); Lewis et al. (2020); Asai et al. (2024); Yan et al. (2024) has emerged as a promising approach, leveraging external knowledge from large-scale databases that integrate an extensive set of sources Vu et al. (2024); Kasai et al. (2024).
@@
Although external databases provide valuable information, they also risk retrieving incorrect information from unreliable sources Pan et al. (2023); Chen et al. (2024); Greshake et al. (2023). This vulnerability stems from a fundamental limitation of retrieval, which relies solely on relevance measures between queries and documents Robertson and Walker (1994); Karpukhin et al. (2020); Ni et al. (2022); Izacard et al. (2022), overlooking source reliability heterogeneity. Furthermore, malicious sources can exploit this limitation by crafting highly relevant yet incorrect documents, leading to misleading outputs Zhong et al. (2023); Zou et al. (2024). While existing methods Weller et al. (2024); Xiang et al. (2024); Deng et al. (2024); Pan et al. (2024) attempt to mitigate this issue by refining retrieved documents, they do not address the retrieval problem itself, allowing unreliable sources to dominate the retrieval process.
@@
To overcome these challenges, we propose Reliability-Aware RAG (RA-RAG), a new multi-source RAG framework that estimates source reliability and effectively integrates it into both the retrieval and aggregation processes. Compared to standard RAG, which retrieves documents without distinguishing between sources, RA-RAG performs source-level retrieval and aggregates information based on estimated reliability using weighted majority voting (WMV), as illustrated in Figure 1. Specifically, RA-RAG consists of two steps. First, given a set of fact-checking queries, we estimate source reliability by cross-checking information across multiple sources without requiring manual fact-checking. This is achieved by leveraging RAG’s ability to automatically retrieve and generate responses (Section 4.2). Second, using the estimated reliability, we propose $\kappa$-reliable and relevant source selection ($\kappa$-RRSS) for WMV, where RA-RAG consults only a small number of reliable sources with relevant documents (Section 4.1). This enhances robustness against unreliable sources while maintaining computational scalability without compromising performance.
The effectiveness of RA-RAG stems from its ability to estimate source reliability, a crucial first step in combating misinformation Popat et al. (2017); Baly et al. (2018, 2020); Burdisso et al. (2024). While source reliability remains underexplored in RAG despite its significance, RA-RAG explicitly incorporates it to improve retrieval and answer generation. Comprehensive experiments and analyses demonstrate that RA-RAG not only effectively estimates source reliability but also robustly aggregates information from multiple sources with heterogeneous reliability. Moreover, it remains scalable even as the number of sources increases. Furthermore, our method effectively estimates the reliability of real-world sources, highlighting its practical applicability. Our main contributions are summarized as follows:
  • We propose RA-RAG, a multi-source RAG framework that estimates source reliability by cross-checking information across multiple sources without relying on manual fact-checking (Section 4.2). Based on the estimated reliability, it retrieves reliable and relevant documents by $\kappa$-RRSS and aggregates them with WMV, generating robust answers while remaining scalable to a large number of sources (Section 4.1).
@@ 2 Related Works
Robust RAG against misinformation. In response to misinformation risks in RAG, several robust methods have been proposed, primarily focusing on improving answer generation after retrieval. Weller et al. (2024); Xiang et al. (2024) utilize majority voting, which is effective only when most retrieved documents are trustworthy. Deng et al. (2024) evaluates document credibility using LLMs’ internal knowledge, but this approach is inherently limited as it misaligns with RAG’s core rationale of leveraging external knowledge to address LLMs’ limitations. Pan et al. (2024) assigns binary credibility scores (high/low) to retrieved documents based on source reputation and incorporates them into prompts. However, this approach is unsuitable for sources with obscure reputations, and reputation does not necessarily reflect actual reliability. In contrast, RA-RAG integrates source reliability into the retrieval and answer generation process by explicitly estimating it. Additional discussion on related works is presented in Appendix D.
@@ 3 Problem Formulation
In this section, we first introduce the standard RAG framework in Section 3.1, widely used in previous works but has a clear limitation: overlooking the source reliability heterogeneity. To address this, we introduce a multi-source RAG framework that accounts for source reliability in Section 3.2, followed by a discussion of its key challenges.
@@ 3.2 Multi-source RAG with source reliability
We introduce a multi-source RAG framework that explicitly distinguishes between the sources of documents and incorporates source reliability. Let $N$ be the number of distinct sources contributing to the database $\mathcal{D}$. We partition the database as $\mathcal{D}=\bigcup_{i=1}^{N}\mathcal{S}_{i}$, where $\mathcal{S}_{i}$ is the set of documents originating from source $i\in[N]$. Such a partition of dataset $\mathcal{D}$ enables the system to account for the reliability of each document’s source, based on weighted majority voting (WMV). For a given query $q$, let $\hat{y}_{i}=\mathcal{G}(q,\mathcal{R}(q,\mathcal{S}_{i}))$ represent the generated response using retrieved documents exclusively from source $S_{i}$. Once the probability of a retrieved document from source $i$ being correct is estimated as $v_{i}$, and a set of candidate responses $\mathcal{M}$ is obtained from $\hat{y}_{i}$’s, we apply WMV to aggregate the responses as follows:
$\hat{y}=\operatorname*{arg\,max}_{u\in\mathcal{M}}\sum_{i\in[N]}v_{i}\mathbbm{1}(\hat{y}_{i}=u)\;.$ (1)
@@
Inherent issues with LLM. LLMs may generate hallucinations or misaligned answers influenced by their internal knowledge Kaddour et al. (2023); Ji et al. (2023); Kortukov et al. (2024); Xu et al. (2024), distorting their alignment with retrieved documents and complicating the WMV process. Additionally, LLMs often generate semantically identical responses with paraphrasing, making response aggregation of $\hat{y}_{i}$’s more challenging.
@@ 4 Method: RA-RAG
We propose Reliability-Aware RAG (RA-RAG) to address key challenges in multi-source RAG. Prior to deployment, RA-RAG estimates source reliability using an iterative reliability estimation algorithm, which cross-checks information across multiple sources through fact-checking queries designed to verify documents within each source (Section 4.2). Leveraging RAG’s ability to retrieve relevant documents and generate responses automatically, RA-RAG enables automated reliability estimation without manual fact-checking. During inference, RA-RAG then aggregates responses from different sources based on the estimated reliability (Section 4.1).
For ease of presentation, we first introduce the aggregation process in Section 4.1, then propose the iterative reliability estimation method in Section 4.2.
@@ 4.1 Aggregation process
Although the instruction prompt guides the model to output “I don’t know" (IDK) when there is no relevant information, LLMs may still produce misaligned responses, undermining effective aggregation. To address this, a filtering function $f_{\text{align}}$ is necessary to detect and replace misaligned responses with IDK. In this work, we utilize AlignScore Zha et al. (2023), which evaluates the factual consistency of a response $\hat{y}_{i}$ relative to the query $q$ and retrieved documents $\mathcal{R}(q,\mathcal{S}_{i})$:
$f_{\text{align}}(\hat{y}_{i},q,\mathcal{R}(q,\mathcal{S}_{i}))$
=$=\begin{cases}\text{IDK}&\text{if }\mathcal{E}(\hat{y}_{i};q,\mathcal{R}(q,\mathcal{S}_{i}))<\tau\\ \hat{y}_{i}&\text{otherwise}\end{cases}\;,$
where $\mathcal{E}$ represents AlignScore function and $\tau$ is threshold. For simplicity, we omit $\mathcal{E}$ in $f_{\text{align}}(\hat{y}_{i},q,\mathcal{R}(q,\mathcal{S}_{i}))$. Further details on the filtering method and threshold are provided in Appendix A. By applying this filtering method, we obtain a refined set of candidate responses:
@@
Additionally, since LLMs often paraphrase responses with equivalent meanings (e.g., “There are 24 hours in a day." vs. “Each day has 24 hours."), we cluster responses in $\mathcal{M}_{\text{filtered}}$ based on semantic equivalence. We denote the refined set obtained through semantic clustering as $\mathcal{C}(\mathcal{M}_{\text{filtered}})=\{C_{k}\subseteq\mathcal{M}_{\text{filtered}}\}_{k=1}^{K}$, where each $C_{k}$ represents a distinct cluster such that $C_{i}\bigcap C_{j}=\emptyset$ for all i $\neq$ j. For semantic clustering method $\mathcal{C}$, we employ the algorithm by Kuhn et al. (2023), which clusters responses that mutually entail each other using a pretrained natural language inference (NLI) model. Following Kuhn et al. (2023), we use the DeBERTa-large model He et al. (2021) for clustering.
@@
$\hat{y}\!=\!\!\!\operatorname*{arg\,max}_{u\in\mathcal{C}(\mathcal{M}_{\text{filtered}})}\!\sum_{i\in[N]}\!v_{i}\mathbbm{1}(f_{\text{align}}(\hat{y}_{i},q,\mathcal{R}(q,\mathcal{S}_{i}))\!=u)\!\;.$ (2)
To generate the final response $\hat{y}$, we select the first response in the cluster $C_{k}$, as all responses within the cluster are considered semantically equivalent.
Efficient aggregation. In real-world applications, aggregating information from all sources can be computationally expensive, especially when the number of sources is large. To mitigate this, we propose $\kappa$-Reliable and Relevant Source Selection ($\kappa$-RRSS). This method iterates over sources in descending order of reliability $v_{i}$ and selects the first $\kappa$ sources that contain relevant information, where $\kappa<N$. A source is deemed irrelevant if its filtered response $f_{\text{align}}(\hat{y}_{i},q,\mathcal{R}(q,\mathcal{S}_{i}))$ is IDK. For the formal algorithm, please refer to Algorithm 1. Given the set of responses from the selected sources, denoted as $\mathcal{M}_{\kappa}$, the final response is aggregated as follows:
$\hat{y}\!=\!\!\!\!\!\operatorname*{arg\,max}_{u\in\mathcal{C}(\mathcal{M}_{\kappa\text{-filtered}})}\!\sum_{i\in[N]}\!\!v_{i}\mathbbm{1}(f_{\text{align}}(\hat{y}_{i},q,\mathcal{R}(q,\mathcal{S}_{i}))\!=u)\!\;,$ (3)
where $\mathcal{M}_{\kappa\text{-filtered}}$ denotes the set of responses from $\mathcal{M}_{\kappa}$ after applying $f_{\text{align}}$. By focusing on reliable and relevant sources, $\kappa$-RRSS significantly reduces inference overhead while maintaining robust performance.
@@ 4.2 Iterative reliability estimation
To estimate source reliability and effectively aggregate outputs, we utilize the WMV method proposed by Li and Yu (2014), a simple yet effective approach for aggregating crowdsourced labels in classification tasks. Specifically, we first generate fact-checking queries for documents. For example, if a document in a source states, “COVID-19 is caused by 5G networks", we can generate a query such as “What causes COVID-19?". Given a set of $M$ fact-checking queries, denoted as $\{q^{j}\mid j\in[M]\}$, the iterative reliability estimation process is described as follows: